OpenStudy (anonymous):

Solve the equation on the interval 0≤theta<2π. (4cos^2)x - 3 ( I'll rewrite this so it makes more sense)

OpenStudy (anonymous):

Solve the equation on the interval \[0\le \theta <2\pi \]Here's the equation:\[4\cos ^{2}x-3=0\]

hartnn (hartnn):

try to isolate cos^2 x from there first, can you ?

OpenStudy (anonymous):

cos^2x=3/4?

OpenStudy (anonymous):

And welcome back (:

hartnn (hartnn):

hey, thanks :) ok, so cos^2x=3/4 take square root on both sides, cos x= ... , ..... ?

OpenStudy (anonymous):

(√3/4)? and yw (:

hartnn (hartnn):

did you mean \(\large \sqrt{\dfrac{3}{4}}\) ?

OpenStudy (anonymous):

Yup o:

hartnn (hartnn):

also when you take square root , you get 2 values of cos x , you know what the other value is ?

hartnn (hartnn):

hint : \(if x^2=a, \implies x=\pm \sqrt{a}\)

OpenStudy (anonymous):

I think you lost me o;

hartnn (hartnn):

\(\cos^2 x =3/4 \\ \implies \cos x = \pm \sqrt{3/4}= \sqrt 3/2, -\sqrt 3/2\) got this ?

OpenStudy (anonymous):

How'd you get ((√3)/2) from √3/4)?

hartnn (hartnn):

\(\sqrt{3/4}= \sqrt3/\sqrt4\) abd square root of 4 is 2 =\(\sqrt 3/2\)

hartnn (hartnn):

*and

OpenStudy (anonymous):

Oh, okay! Go on (:

hartnn (hartnn):

for what values of x in unit circle 0≤theta<2π. is cos x = sqrt 3/2 ?? for what values of x in unit circle 0≤theta<2π. is cos x = -sqrt 3/2 ??

OpenStudy (anonymous):

π/6, 5π/6, 7π/6, 11π/6 (:

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

\(\huge \checkmark\) welcome ^_^